OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each f(x)=5(x+7)^2(x-7)^3

OpenStudy (anonymous):

can you walk me through the steps @***[ISURU]*** ?

OpenStudy (anonymous):

@***[ISURU]***

OpenStudy (***[isuru]***):

How far did you went on this one ?... did you manage to find the zeros of this polynomial ?

OpenStudy (anonymous):

I do not know how to work this problem. I need someone to show me how to solve it. :D

OpenStudy (***[isuru]***):

lets take the polynomial p(x) = 5(x+7)(x-7) Its the same polynomial as your question but without the powers of 2 and 3 now to find the zeros... lets find the "x" values when p(x) = 0 0 = 5(x+7)(x -7) now can you find the "x" values where this function will become zero ?

OpenStudy (anonymous):

would you do x+7=0 and x-7=0 ?

OpenStudy (***[isuru]***):

yes...and what will you get ? :)

OpenStudy (anonymous):

-7 and 7

OpenStudy (anonymous):

ok so the answer would -7 multiplicity 2 and 7 multiplicity 3 right?

OpenStudy (***[isuru]***):

You got it correct ! :)

OpenStudy (anonymous):

thank you!

OpenStudy (***[isuru]***):

you r welcome :)